Carbon monoxide reacts with oxygen gas to form CO2, as shown by the balanced equation below.

2CO (g) + O2 (g) ----------> 2CO2 (g)

A reaction vessel contains 3.000 g of CO and 3.000 g of O2.

What is the maximum number of moles of CO2 recovered, based on the number of moles of the given reactants?
Answer #1

2 CO(g) + O2(g) --------------- 2 CO2(g)

2 mole     1 mole                      2 mole

mass of CO = 3.000 g

molar mass of CO = 28.0 gram/mole

number of moles of CO = mass/molar mass = 3.000/28.0 = 0.107mole

mass of O2 = 3.000 g

molar mass of O2 = 32.0 gram/mole

number of moles of O2 = 3.000/32.0 = 0.09375 moles

according to balanced equation

1 mole of O2 = 2 moles of CO

0.09375 moles of O2 = ?

                                     = 2 x 0.09375 /1 = 0.1875 moles of CO

we need 0.1875 moles of CO. but we have 0.107 moles of CO. So, CO is consumed first in the reaction.

Hence CO is limiting reagent

according to equation

2 moles of CO = 2 moles of CO2

0.107 moles of CO = ?

                                  = 2 x 0.107/2 = 0.107 moles of CO2

number of moles of CO2 formed = 0.107 moles.